OpenStudy (anonymous):

What is the equation of the line passing through (2, -3) and parallel to the line y = -6x - 1 in standard form? 6x + y = 9 6x - y = -9 x + 6y = 9 x - 6y = -9

OpenStudy (anonymous):

Do you know what parallel means?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

OK, what does it mean and what does it mean we have to do with out eqaution?

OpenStudy (anonymous):

the lines are side by side

OpenStudy (anonymous):

do the equations have to be similar or

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The only similar part is their slope

OpenStudy (anonymous):

The slopes are the same with parallel lines so that's why they never touch

OpenStudy (anonymous):

Now you have to find the slope of y = -6x - 1, our given equation and that will give us the slope of our parallel line (since parallel lines have the same slope)

OpenStudy (anonymous):

helpppppp:( idk howwww

OpenStudy (anonymous):

isnt it just -6????

OpenStudy (anonymous):

Yes, it's the number on the x

OpenStudy (anonymous):

isnt it A

OpenStudy (anonymous):

Now we have our slope of the parallel line (-6) and they say it passes through the point (2,-3), so we use our handy-dandy formula and plug in \[\Huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) is a point

OpenStudy (anonymous):

\[\Large (y-(-3))=-6(x-2)=~~~~y+3=-6x+12~~~~~~~~6x+y=9\]

OpenStudy (anonymous):

6x+y=9

OpenStudy (anonymous):

youre the bomb

OpenStudy (anonymous):

So A is correct, and those are your steps!

OpenStudy (anonymous):

And yw